OpenStudy (anonymous):

How can an expression or process be determined for an arithmetic sequence? medals will be given :)

OpenStudy (matlee):

Lols i havve flvs too what school do you go to?

OpenStudy (anonymous):

lol my zoned school?

OpenStudy (matlee):

yeah

OpenStudy (matlee):

my school ends in 2 weeks

OpenStudy (anonymous):

oh lol. and lake nona high wbu

OpenStudy (matlee):

east river

OpenStudy (matlee):

your soccer team is hard lol :l

OpenStudy (anonymous):

LMAO

OpenStudy (anonymous):

east river is like an hour away from me

OpenStudy (matlee):

I shall beat them next year ok what do you need help with?

OpenStudy (matlee):

Oh ok , so just write an expalnation?

OpenStudy (anonymous):

lol. and i just need help with the question at the top

OpenStudy (anonymous):

yep

OpenStudy (matlee):

@Here_to_Help15 hey can you help us :o

OpenStudy (anonymous):

what grade are you in

OpenStudy (matlee):

11th

OpenStudy (matlee):

u?

OpenStudy (matlee):

12th next year

OpenStudy (anonymous):

10th ._.

OpenStudy (matlee):

:O

OpenStudy (anonymous):

lmfao oh gosh

OpenStudy (anonymous):

@Here_to_Help15 you able to help?

OpenStudy (matlee):

@Nnesha Hi Nnesha ! If your nt busy we would really like your help!

OpenStudy (anonymous):

i appreciate you finding help for me lol

OpenStudy (matlee):

Np! Lol im getting off after this tahts why :p you still have exams?

OpenStudy (anonymous):

like eoc's? nah

OpenStudy (matlee):

Oh thats right this is for flvs lol

OpenStudy (anonymous):

haha

OpenStudy (here_to_help15):

Is this is a write out answer or multiple choice?

OpenStudy (anonymous):

write out answer

OpenStudy (here_to_help15):

The @Nnesha will be able to help :)

Nnesha (nnesha):

arithmetic sequence: adding same number from one term to to the next formula for arithmetic seq \[\huge\rm a_n = a_1 +(n-1)d\] where a_1= first term d= common difference

Nnesha (nnesha):

the number would be same every time for example 2,4,6,8 common difference is 2 2+2=4 4+2=6 6+2=8

OpenStudy (anonymous):

thank you so much you guys

Nnesha (nnesha):

my pleasure :-)